# sources.bib

### v1

@@ 2 Superspaces
A family of supermanifolds is a flat submersion of supermanifolds $M\to A$. We say that $M$ has relative dimension $\dim(M)-\dim(A)$.
Note that since we defined supermanifolds as locally ringed spaces, they technically do not possess isolated geometrical points: In fact, we can only look at the stalks, which capture infinitesimal neighbourhoods of points. However, upon considering families of supermanifolds $M\to A$, we can consider $A$-points of $M$, which are sections of the submersion $A\to M$. By viewing all supermanifolds over $A$ in this way, we can use the odd directions of $A$ to mix even and odd components of $M$ and $N$ in morphisms $f:M\to N$.
@@
We can replace the base by pulling back along a morphism of supermanifolds $\tilde{A}\to A$, giving rise to a contravariant functor
$\mathbf{sMfd}_{A}\to\mathbf{sMfd}_{\tilde{A}}.$ (2)
@@
As shown in Ref. [kessler2018supergeometry], this geometrical implementation of the functor-of-points allows for notions of tangent bundles, super Lie groups, principal fibre bundles, principal connections and many more. In particular, we shall use the principal bundle of frames and replacements of its structure group [^1] The misnomer reductions of the structure group is usually used; but the supporting group homomorphisms need not be injective in general, such that we resort to the use of this terminology. to express super gauge theories geometrically.
@@
$F_{\mathcal{F}}:\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{B},\quad F_{\mathcal{F}}(X\wedge Y)=\operatorname{pr}_{\mathcal{B}}([X,Y])$ (4)
@@
To give physical context, elements of $\mathcal{B}\subset\mathcal{T}_{M}$ will be thought of as bosonic fields, and elements of $\mathcal{F}\subset\mathcal{T}_{M}$ will be thought of as fermionic (spinorial) fields. We shall call $B$ the bosonic distribution and $\mathcal{F}$ the spinorial distribution. The Frobenius map is the obstruction to $\mathcal{F}$ being integrable, and can be thought of as a current:
@@
$Q^{\mathcal{B}}\times_{M}Q^{\mathcal{F}}\to\operatorname{Fr}(TM)\to M.$ (6)
Here, we consider $\operatorname{GL}(m|n)$ as the super Lie group of even automorphisms of $\mathbb{R}^{m|n}$, viewed as a trivial family of supermanifolds over the auxiliary base $A$. The reduction decomposes the soldering form
$\theta_{(p_{\mathcal{B}},p_{\mathcal{F}})}=(p_{\mathcal{B}},p_{\mathcal{F}})^{-1}\circ\mathrm{d}\pi=p_{\mathcal{B}}^{-1}\circ\operatorname{pr}_{\mathcal{B}}^{*}\mathrm{d}\pi+p_{\mathcal{F}}^{-1}\circ\operatorname{pr}_{\mathcal{F}}^{*}\mathrm{d}\pi=\theta_{p_{\mathcal{B}}}^{\mathcal{B}}+\theta_{p_{\mathcal{F}}}^{\mathcal{F}}.$ (7)
@@
Let $\varphi$ be a $\operatorname{GL}(m)\times\operatorname{GL}(n)$ connection on the superspace, let $\Theta^{\varphi}$ denote its torsion, and let $X,Y$ be horizontal lifts of $\tilde{X},\tilde{Y}\in\mathcal{F}\subset\mathcal{T}_{M}$ with respect to $\varphi$. It holds
$\operatorname{pr}_{V}\left(\Theta^{\varphi}(X,Y)\right)=\iota_{[X,Y]}\theta^{\mathcal{B}}.$ (8)
@@
$\operatorname{pr}_{V}\Theta^{\varphi}(X,Y)=\iota_{X}\iota_{Y}\mathrm{d}^{\varphi}\theta^{\mathcal{B}}=\iota_{X}L_{Y}\theta^{\mathcal{B}}=\iota_{[X,Y]}\theta^{\mathcal{B}},$ (9)
where we used that $\theta^{\mathcal{B}}$ vanishes on $\operatorname{Spin}(V,q)$-invariant lifts of elements of $\mathcal{F}$. ∎
@@ 3 Spin Structures
Throughout, we shall denote by $(V,q)$ the quadratic vector space that infinitesimally models the ordinary spacetime manifold, and by $S$ a real spinorial representation of $\operatorname{Spin}(V,q)$. We assume throughout that $(V,q)$ is Minkowskian. The following theorems provide existence and uniqueness results of the $\Gamma$ maps.
@@
Let $S$ be an irreducible real spinorial representation of $(V,q)$, where $q$ is Minkowskian. There exists, up to rescaling, a unique real symmetric morphism of representations $\Gamma:S\otimes S\to V$. It holds
  • $\Gamma$ is invariant under the diagonal action of $K\subset\mathbb{F}$.
  • Let $v\in V$ such that $q(v,v)>0$. The bilinear form $q(v,\Gamma(\cdot,\cdot))$ is non-degenerate.
@@
Let $S$ be a real spinorial representation (not necessarily irreducible), and $\Gamma:S\otimes S\to V$ a symmetric map of representations of $\operatorname{Spin}(V,q)$. It holds:
@@
  • If each of the $\Gamma^{(i)}$ is non-zero, there exists a unique $\tilde{\Gamma}:S^{\vee}\otimes S^{\vee}\to V$ such that the dualisations $\gamma:V\to\operatorname{Hom}(S,S^{\vee})$, $\tilde{\gamma}:V\to\operatorname{Hom}(S^{\vee},S)$ (after employing the musical isomorphism due to $q$ on $V$) obey the Clifford relations
@@
These algebraic relations capture infinitesimal displacements of fermionic and bosonic particles. In what follows, we define the global extension of these relations by imposing distinguished spin structures on superspaces.
@@
  • the bosonic distribution $\mathcal{B}$ has typical fibre $V$,
  • the spinorial distribution $\mathcal{F}$ has typical fibre $\Pi S$, where $S$ is a real spinorial representation of $\operatorname{Spin}(V,q)$, and
@@
    $Q^{\mathcal{F}}\to Q^{\mathcal{B}}\times_{M}Q^{\mathcal{F}}\to\operatorname{Fr}(TM)$ (13)
@@
    $\operatorname{Spin}(V,q)\xrightarrow{\alpha\times\operatorname{id}}\operatorname{SO}(V,q)\times\operatorname{Spin}(V,q)\to\operatorname{GL}(m|n).$ (14)
For brevity, we shall write $Q^{\mathcal{F}}\eqqcolon Q$.
In particular, this implies that the map $\operatorname{pr}_{\mathcal{B}}([\cdot,\cdot])$ respects the reduction, such that by using the soldering form, it can be viewed as a map
$\operatorname{pr}_{\mathcal{B}}([\cdot,\cdot])\in\Omega^{0|2}(Q,V)^{\operatorname{Spin}(V,q)}_{\operatorname{hor}}.$ (15)
@@
The bosonic distribution is integrable, that is, there exists a purely even submanifold $\check{M}\subset M$ such that $\mathcal{B}|_{\check{M}}=\mathcal{T}_{\check{M}}$.
@@
We need to show that the Frobenius obstruction $\operatorname{pr}_{\mathcal{F}}([\cdot,\cdot]|_{\mathcal{B}\wedge\mathcal{B}})$ vanishes. To that end, note that the obstruction presented in this way is a graded skew $\mathcal{O}_{M}$-bilinear map $\mathcal{B}\wedge\mathcal{B}\to\mathcal{F}$. By means of the soldering form, it can be lifted to a bilinear $\operatorname{Spin}(V,q)$-equivariant map $V\otimes V\to S$. Suppose the map was non-vanishing for some $v\otimes w\in V\otimes V$. But $-1\in\operatorname{Spin}(V,q)$ is in the kernel of the action on $V\otimes V$, whereas it changes the sign on $S$ – a contradiction. Hence, the Frobenius obstruction vanishes. ∎
@@
If the almost spinorial superspace is of type $\mathcal{N}=1$, the Frobenius map coincides with the morphism induced by the map of representations $\Gamma:S\otimes S\to V$.
@@
Note first that the Frobenius pairing is indeed $\mathcal{O}_{M}$-bilinear, such that by the associated bundle construction and by using the soldering form, it lifts to a $\operatorname{Spin}(V,q)$-equivariant map $\operatorname{Sym}^{2}(S)\to V$, which is non-degenerate. By theorem 1, these maps are uniquely (up to rescaling) given by the maps of representations from theorem 1. ∎
@@
An almost spinorial superspace is called spinorial superspace if the Frobenius obstruction $\operatorname{pr}_{\mathcal{B}}([\cdot,\cdot])$ coincides with the map induced by the morphism of $\operatorname{Spin}(V,q)$ representations $\Gamma:S\otimes S\to V$.
@@
Henceforth, we shall exclusively consider spinorial superspaces, and embrace the fact that spinorial superspaces encode their spinorial nature in the spinorial distribution $\mathcal{F}\subset\mathcal{T}_{M}$.
@@
Furthermore, the map $\Gamma:\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{B}$ is non-degenerate in the sense that for non-isotropic $V\in\mathcal{B}$, the bilinear form $g(\Gamma(\cdot,\cdot),V)$ is non-degenerate.
@@
In particular, the $\mathcal{O}_{M}$-bilinear map $\operatorname{pr}_{\mathcal{B}}([\cdot,\cdot])$ respects the reduction, such that it can again be viewed as the $\Gamma$ map
@@
Every spinorial superspace can be turned into a $\operatorname{Spin}^{\mathbb{F}}(V,q)$ superspace, by means of the central product
@@
To that end, note that we have a morphism of principal bundles $Q\to Q^{\mathcal{B}}$, along the double covering $A:\operatorname{Spin}(V,q)\to\operatorname{SO}(V,q)$. Since the covering is discrete, $A_{*}:\mathfrak{spin}(V,q)\to\mathfrak{so}(V,q)$ is an isomorphism of Lie algebras. Thus, connections on $Q^{\mathcal{B}}$ isomorphically lift to connections on $Q$. Note that $\mathcal{A}_{Q^{\mathcal{B}}}$ is affine over $\Omega^{1}(Q^{\mathcal{B}},\mathfrak{so}(V,q))^{\operatorname{SO}(V,q)}_{\operatorname{hor}}$. Recall that the torsion of a connection can be interpreted via the exact sequence of $\mathcal{O}_{M}$-modules
@@
where $\hat{\alpha}(a)=\rho_{*}(a)\wedge\theta_{B}$, and $\rho_{*}:\mathfrak{so}(V,q)\to\operatorname{End}(V)$ is the induced map of Lie algebras from the $\operatorname{SO}(V,q)$-structure $Q^{\mathcal{B}}$. $\hat{K}$ and $\hat{C}$ denote kernel and cokernel of $\hat{\alpha}$, respectively. This sequence arises from vector bundles associated to $Q^{\mathcal{B}}$ via
@@
with subsequent antisymmetrisation on the first two factors. Clearly, $\alpha$ is an isomorphism (and in particular, $\hat{K}=\hat{C}=\{0\}$). Thus, so is $\hat{\alpha}$; and since $\mathcal{A}_{Q^{\mathcal{B}}}$ is affine over $\Omega^{1}(Q^{\mathcal{B}},\mathfrak{so}(V,q))^{\operatorname{SO}(V,q)}_{\operatorname{hor}}$, the map
$\mathcal{A}_{Q^{B}}\to\Omega^{2}(Q^{\mathcal{B}},V)^{\operatorname{SO}(V,q)}_{\operatorname{hor}},\quad\tilde{\varphi}\mapsto\mathrm{d}^{\tilde{\varphi}}\theta_{\mathcal{B}}$ (32)
@@
$\mathcal{A}^{s}_{Q}\cong\mathcal{A}_{Q^{K}}$ (34)

### v2

@@ 2 Superspaces
A family of supermanifolds is a surjective submersion of supermanifolds $M\to A$. We say that $M$ has relative dimension $\dim(M)-\dim(A)$.
Note that since we defined supermanifolds as locally ringed spaces, they technically do not possess isolated geometrical points: In fact, we can only look at the stalks, which capture infinitesimal neighbourhoods of points. However, upon considering families of supermanifolds $M\to A$, we can consider $A$-points of $M$, which are sections $A\to M$ of the submersion. By viewing all supermanifolds over $A$ in this way, we can use the odd directions of $A$ to mix even and odd components of $M$ and $N$ in morphisms $f:M\to N$. In other words, we use the auxiliary base space $A$ to parametrise the supermanifold $M$.
@@
We can replace the base by pulling back the submersion $\sigma:M\to A$ along a surjective submersion of supermanifolds $\varphi:A\to\tilde{A}$, giving rise to a covariant functor
$\mathbf{sMfd}_{A}\to\mathbf{sMfd}_{\tilde{A}},\quad(M,\sigma,A)\mapsto(M,\sigma\circ\varphi,\tilde{A}).$ (2)
@@
As shown in Ref. [kessler2018supergeometry], this geometrical implementation of the functor-of-points allows for notions of tangent bundles, flows of vector fields, super Lie groups, principal fibre bundles, principal connections and many more. In particular, we shall use the principal bundle of frames and replacements of its structure group [^1] The misnomer reductions of the structure group is usually used; but the supporting group homomorphisms need not be injective in general, such that we resort to the use of this terminology. to express super gauge theories geometrically.
We shall now give a enlightening example as to why the auxiliary base is essential to make sense of some geometrical constructions. Recall that the flow of a vector field $X$ is given by the family of diffeomorphisms $\phi^{X}_{t}:M\to M$ such that
$X(f)\circ\phi_{t}^{X}=\frac{\mathrm{d}}{\mathrm{d}t}f\circ\phi_{t}^{X}.$ (3)
Example 1.
Let $M$ be a supermanifold, and denote by $U\subset M$ a chart domain. In what follows, we consider flows of local coordinate vector fields. To that end, let $(x^{a},\xi^{\alpha})$ denote coordinate functions on $U$. Let us denote by $\mathcal{O}_{U}$ the restriction of $\mathcal{O}_{M}$ to $U$, and similarly, write $\mathcal{T}_{U}$ for the restricted tangent sheaf. Note that
$\mathcal{T}_{U}=\operatorname{span}_{\mathcal{O}_{U}}\left\{\frac{\partial}{\partial x^{a}},\frac{\partial}{\partial\xi^{\alpha}}\right\}=\operatorname{span}_{\mathcal{O}_{U}}\{\frac{\partial}{\partial X_{A}}\},$ (4)
where
$X_{A}=\begin{cases}{x^{a}}&\textrm{if }A=a,\\ {\xi^{\alpha}}&\textrm{if }A=\alpha.\end{cases}$ (5)
An explicit computation shows that in these coordinates, the flow $\phi^{\frac{\partial}{\partial X^{B}}}_{t}\eqqcolon\phi^{B}_{t}$ is given by
$\phi^{B}_{t}:U\to U,\quad(X^{1},\ldots,X^{B-1},X^{B},X^{B+1},\ldots)\mapsto(X^{1},\ldots,X^{B-1},X^{B}+t,X^{B+1},\ldots),$ (6)
that is, for short, $\phi^{B}_{t}:X^{A}\mapsto X^{A}+t\delta^{AB}$. In particular, whenever $X^{B}$ is an odd coordinate, $t$ needs to be an odd parameter. Thus, instead of taking $\mathbb{R}$ as ground ring, we require some graded ring $R$ with non-trivial odd component. The submersion $U\to A$ in the category $\mathbf{sMfd}_{A}$ allows us to take $R=\mathcal{O}_{A}$, giving the ground ring geometric significance.
@@
$\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{B},\quad(X\wedge Y)\mapsto[X,Y]\mod\mathcal{F}$ (8)
@@
Recall that by the Serre-Swan theorem for supermanifolds (see, e.g., Ref. [kessler2018supergeometry]), locally free $\mathcal{O}_{M}$-modules correspond uniquely to (locally free) vector bundles over $M$, such that the section functor
$\mathbf{lfVB}_{M}\xrightarrow{\Gamma(M,-)}\mathbf{lfMod}_{\mathcal{O}_{M}},\quad E\mapsto\Gamma(M,E)\eqqcolon\mathcal{E}$ (9)
defines an equivalence of categories. In this spirit, we shall write $TM\to M$ for the vector bundle that corresponds to the tangent sheaf, and likewise, $F$ and $B$ for the regular distributions that correspond to $\mathcal{F}$ and $\mathcal{B}$, that is,
$\mathcal{B}=\Gamma(M,B),\quad\mathcal{F}=\Gamma(M,F).$ (10)
We will freely choose between the sheaf language and the vector bundle language, depending on which description suits us best.
Let us now elaborate on the second point of this definition. To that end, recall that once we specify a distribution $\mathcal{F}$, we have the $\mathcal{O}_{M}$-bilinear Frobenius map
$\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{T}_{M}/\mathcal{F},\quad X\wedge Y\mapsto[X,Y]\mod\mathcal{F}.$ (11)
This pairing is precisely the obstruction to $\mathcal{F}$ being integrable, i.e., the obstruction to $\mathcal{F}$ arising as a foliation of $M$ by subsupermanifolds $N_{\alpha}$ such that the restricted tangent sheaves $\left.\mathcal{T}_{M}\right|_{N_{\alpha}}$ arise as
$\left.\mathcal{T}_{M}\right|_{N_{\alpha}}\cong\left.\mathcal{O}_{M}\right|_{N_{\alpha}}\otimes\mathcal{T}_{N_{\alpha}}.$ (12)
Note that in our case, we additionally have orthogonal projections $\operatorname{pr}_{\mathcal{F}}:\mathcal{T}_{M}\to\mathcal{F}$ and $\operatorname{pr}_{\mathcal{B}}:\mathcal{T}_{M}\to\mathcal{B}$, such that we can identify the quotient $\mathcal{T}_{M}/\mathcal{F}$ with the orthogonal complement $\mathcal{B}$. This allows us to make the Frobenius obstruction explicit:
$f_{F}:\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{B},\quad X\wedge Y\mapsto\operatorname{pr}_{\mathcal{B}}([X,Y]).$ (13)
To give physical context, elements of $\mathcal{B}\subset\mathcal{T}_{M}$ will be thought of as bosonic fields, and elements of $\mathcal{F}\subset\mathcal{T}_{M}$ will be thought of as fermionic (spinorial) fields. We shall call $\mathcal{B}$ the bosonic distribution and $\mathcal{F}$ the spinorial distribution. The Frobenius map is the obstruction to $\mathcal{F}$ being integrable, and can be thought of as a current:
@@
$Q^{B}\times_{M}Q^{F}\to\operatorname{Fr}(TM)\to M.$ (15)
Here, we consider $\operatorname{GL}(m|n)$ as the super Lie group of even automorphisms of $\mathbb{R}^{m|n}$, viewed as a trivial family of supermanifolds over the auxiliary base $A$. The reduction decomposes the soldering form as follows: Let $p=(p_{B},p_{F})$ be a local section of $(Q^{B}\times_{M}Q^{F})|_{U}\to U\subset M$. Such a section is by definition a local basis of $B\oplus F$, such that it amounts to the respective isomorphisms
$p_{B}:\mathbb{R}^{m|0}\times U\to B|_{U},\quad p_{F}:\mathbb{R}^{0|n}\times U\to F|_{U}.$ (16)
Hence, the reduced soldering form on $Q^{B}\times_{M}Q^{F}$ reads
$\theta_{(p_{B},p_{F})}$ $=(p_{B},p_{F})^{-1}\circ\mathrm{d}\pi=p_{B}^{-1}\circ\operatorname{pr}_{B}^{*}\mathrm{d}\pi\oplus p_{\mathcal{F}}^{-1}\circ\operatorname{pr}_{F}^{*}\mathrm{d}\pi$ (17)
$=\theta_{p_{B}}^{B}\oplus\theta_{p_{F}}^{F}\in\Omega^{1}(Q^{B}\times_{M}Q^{F},\mathbb{R}^{m|0}\oplus\mathbb{R}^{0|n})^{\operatorname{GL}(m)\times\operatorname{GL}(n)}_{\operatorname{hor}}.$ (18)
@@
Let $\varphi$ be a $\operatorname{GL}(m)\times\operatorname{GL}(n)$ connection on the superspace, let $\Theta^{\varphi}$ denote its torsion, and let $X,Y\in\mathcal{T}_{Q^{B}\times_{M}Q^{F}}$ be the horizontal lifts of $\tilde{X},\tilde{Y}\in\mathcal{F}\subset\mathcal{T}_{M}$ with respect to $\varphi$. It holds
$\operatorname{pr}_{\mathbb{R}^{m|0}}\left(\Theta^{\varphi}(X,Y)\right)=\iota_{[X,Y]}\theta^{B},$ (19)
irrespective of the choice of $\varphi$.
@@
$\operatorname{pr}_{\mathbb{R}^{m|0}}\Theta^{\varphi}(X,Y)=\iota_{X}\iota_{Y}\mathrm{d}^{\varphi}\theta^{B}=\iota_{X}L_{Y}\theta^{B}=\iota_{[X,Y]}\theta^{B},$ (20)
where we used that $\theta^{\mathcal{B}}$ vanishes on $\operatorname{GL}(m|0)\times\operatorname{GL}(0|n)$-invariant lifts of elements of $\mathcal{F}$ (see (17)). In particular, the result does not depend on the choice of connection. ∎
@@ 3 Spin Structures
Throughout, we shall denote by $(V,q)$ the quadratic vector space that infinitesimally models the ordinary spacetime manifold, and by $S$ a real spinorial representation of $\operatorname{Spin}(V,q)$. Any irreducible representation $S_{0}$ has a type $\mathbb{F}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$, such that the compact group of unitary elements $K^{\mathbb{F}}=\{x\in\mathbb{F}:|x|^{2}=1\}$ is (isomorphic to) the $\operatorname{Spin}(V,q)$-invariant automorphisms of $S_{0}$. In what follows, we shall assume that $(V,q)$ is Minkowskian. The following theorems provide existence and uniqueness results of the $\Gamma$ maps.
@@
Let $S_{0}$ be an irreducible real spinorial representation of $(V,q)$, where $q$ is Minkowskian. There exists, up to rescaling, a unique real symmetric morphism of representations $\Gamma:S_{0}\otimes S_{0}\to V$. It holds
  • $\Gamma$ is invariant under the action of $K^{\mathbb{F}}\subset\mathbb{F}$ on $S_{0}$.
  • Let $v\in V$ such that $q(v,v)>0$. The bilinear form $q(v,\Gamma(-,-))$ is non-degenerate.
@@
Let $S$ be a real representation of $\operatorname{Spin}(V,q)$ (not necessarily irreducible), and $\Gamma:S\otimes S\to V$ a symmetric map of representations of $\operatorname{Spin}(V,q)$. It holds:
@@
  • If each of the $\Gamma^{(i)}$ is non-zero, there exists a unique $\tilde{\Gamma}:S^{\vee}\otimes S^{\vee}\to V$ such that the duals $\gamma:V\to\operatorname{Hom}(S,S^{\vee})$, $\tilde{\gamma}:V\to\operatorname{Hom}(S^{\vee},S)$ (after employing the musical isomorphism due to $q$ on $V$) obey the Clifford relations
@@
The super Poincaré algebra is given by the semidirect product of Lie algebras $\mathfrak{p}=(V\oplus\Pi S)\rtimes\mathfrak{spin}(V,q)$. The preceding algebraic relations capture infinitesimal displacements of fermionic and bosonic particles. In what follows, we define the global extension of these relations by imposing distinguished spin structures on superspaces.
@@
    $Q^{F}\to Q^{B}\times_{M}Q^{F}\to\operatorname{Fr}(TM)$ (24)
@@
    $\operatorname{Spin}(V,q)\xrightarrow{\alpha\times\operatorname{id}}\operatorname{SO}(V,q)\times\operatorname{Spin}(V,q)\to\operatorname{GL}(m|n),$ (25)
  • the bosonic distribution $B\subset TM$ has typical fibre $V$ such that $B\cong Q^{B}\times_{\operatorname{SO}(V,q)}V$, and
  • the spinorial distribution $F\subset TM$ has typical fibre $\Pi S$, where $S$ is a real spinorial representation of $\operatorname{Spin}(V,q)$, such that $F\cong Q^{F}\times_{\operatorname{Spin}(V,q)}\Pi S$.
For brevity, we shall write $Q^{F}\eqqcolon Q$.
Note that the Frobenius obstructions can be made explicit via
$f_{F}:\wedge^{2}F\xrightarrow{[-,-]}TM\xrightarrow{\operatorname{pr}_{B}}B,$ (26)
as the orthogonal projection onto $B$ identifies $TM/F\cong B$.
The approach to superspaces via the bundle of frames shows that Frobenius obstructions to $\mathcal{B}$ $\mathcal{F}$ must arise as maps of $\operatorname{Spin}(V,q)$ representations. We illustrate this for the integrability conditions of $\mathcal{F}$. To do so, we pick any local section $p$ of $Q=Q^{F}$ over $U\subset M$. Recall that we have a map $\hat{\alpha}:Q^{F}\to Q^{B}$, such that $p$ gives rise to the following isomorphisms
$p:\Pi S\times U\xrightarrow{\sim}{}\left.F\right|_{U},\quad\alpha(p):V\times U\xrightarrow{\sim}{}\left.B\right|_{U}.$ (27)
Using the inverse of the maps induced by the section, we see that $f_{F}$ descends fibrewise to a map of $\operatorname{Spin}(V,q)$ representations, as the following diagram commutes by construction:
[…]
As all local sections $p$ are related by a gauge transformation $g\in C^{\infty}(U,\operatorname{Spin}(V,q))$ and the preceding result is independent of $g$, it follows that the map $\hat{\alpha}(p)^{-1}(f_{F})$ is $\operatorname{Spin}(V,q)$-equivariant, and can thus be identified with an element
$C_{F}\in\Omega^{2}(Q,V)^{\operatorname{Spin}(V,q)}_{\operatorname{hor}}.$ (29)
Note that equivalently, $C_{F}$ can be constructed by using the hom-tensor adjunction
$\operatorname{Hom}(\wedge^{2}(\mathcal{F}),\mathcal{B})\cong\wedge^{2}(\mathcal{F})^{\vee}\otimes_{\mathcal{O}_{M}}\mathcal{B}.$ (30)
Alternatively, using a $\operatorname{Spin}(V,q)$ connection $\varphi\in\mathcal{A}(Q)$, one can lift sections of $\mathcal{F}$ horizontally to $\mathcal{T}_{Q}$ and use the $\operatorname{Spin}(V,q)$-equivariance of the lifted sections.
@@
The bosonic distribution is integrable along any purely even submanifold $\iota:\check{M}\to M$ whenever $\underline{\check{M}}=\underline{M}$ and $\underline{\iota}=\operatorname{id}_{\underline{M}}$.
@@
Note first that $\mathrm{d}\iota:T\check{M}\to\iota^{*}TM$ corestricts to an isomorphism
$T\check{M}\cong\iota^{*}B\cong B_{\iota(\check{M})}.$ (31)
For integrability, it suffices to show that the Frobenius obstruction $f_{B}\coloneqq\operatorname{pr}_{F}([-,-]|_{\mathcal{B}\wedge\mathcal{B}})$ vanishes. To that end, note that the obstruction presented in this way is a graded skew $\mathcal{O}_{M}$-bilinear map $\mathcal{B}\wedge\mathcal{B}\to\mathcal{F}$. As illustrated above, $f_{B}$ corresponds fibrewise to a bilinear $\operatorname{Spin}(V,q)$-equivariant map $V\otimes V\to S$. Suppose the map was non-vanishing for some $v\otimes w\in V\otimes V$. Then, $-1\in\operatorname{Spin}(V,q)$ is in the kernel of the action on $V\otimes V$, whereas it changes the sign on $S$ – a contradiction. Hence, the Frobenius obstruction vanishes. ∎
@@
If the almost spinorial superspace is of type $\mathcal{N}=1$, the Frobenius map $C_{F}$ from equation (29) coincides with the morphism induced by the map of representations $\Gamma:S\otimes S\to V$.
@@
Note first that the Frobenius pairing is indeed $\mathcal{O}_{M}$-bilinear, such that by the associated bundle construction and by using the soldering form, it lifts to a $\operatorname{Spin}(V,q)$-equivariant map $\operatorname{Sym}^{2}(S)\to V$, which is non-degenerate. These maps are unique (up to rescaling), and described by theorem 1. ∎
@@
An almost spinorial superspace is called spinorial superspace if the Frobenius map $C_{F}$ coincides with the map induced by the morphism of $\operatorname{Spin}(V,q)$ representations $\Gamma:S\otimes S\to V$ from theorem 2.
@@
Henceforth, we shall exclusively consider spinorial superspaces, and embrace the fact that spinorial superspaces encode their spinorial nature in the spinorial distribution $\mathcal{F}\subset\mathcal{T}_{M}$. Furthermore, inspired by physics literature, we will also use $\Gamma$ to refer to the Frobenius map $C_{F}$.
@@
Furthermore, the map $\Gamma:\wedge^{2}_{\mathcal{O}_{M}}\mathcal{F}\to\mathcal{B}$ is non-degenerate in the sense that for non-isotropic $v\in\mathcal{B}$, the bilinear form $g(\Gamma(-,-),v)$ is non-degenerate.
@@
In particular, the $\mathcal{O}_{M}$-bilinear map $\operatorname{pr}_{B}([-,-])$ respects the reduction, such that it can again be viewed as the $\Gamma$ map
@@
Every spinorial superspace can be turned into a $\operatorname{Spin}^{\mathbb{F}}(V,q)$ superspace, by setting
@@
To that end, note that we have a morphism of principal bundles $Q\to Q^{B}$, along the double covering $A:\operatorname{Spin}(V,q)\to\operatorname{SO}(V,q)$. Since the covering is discrete, $A_{*}:\mathfrak{spin}(V,q)\to\mathfrak{so}(V,q)$ is an isomorphism of Lie algebras. Thus, connections on $Q^{B}$ isomorphically lift to connections on $Q$. Note that $\mathcal{A}_{Q^{B}}$ is affine over $\Omega^{1}(Q^{B},\mathfrak{so}(V,q))^{\operatorname{SO}(V,q)}_{\operatorname{hor}}$. Recall that the torsion of a connection can be interpreted via the exact sequence of $\mathcal{O}_{M}$-modules
@@
where $\hat{\alpha}(a)=\rho_{*}(a)\wedge\theta_{B}$, and $\rho_{*}:\mathfrak{so}(V,q)\to\operatorname{End}(V)$ is the induced map of Lie algebras from the $\operatorname{SO}(V,q)$-structure $Q^{B}$. $\hat{K}$ and $\hat{C}$ denote kernel and cokernel of $\hat{\alpha}$, respectively. This sequence arises from vector bundles associated to $Q^{B}$ via
@@
with subsequent antisymmetrisation on the first two factors. Clearly, $\alpha$ is an isomorphism (and in particular, $\hat{K}=\hat{C}=\{0\}$). Thus, so is $\hat{\alpha}$; and since $\mathcal{A}_{Q^{B}}$ is affine over $\Omega^{1}(Q^{B},\mathfrak{so}(V,q))^{\operatorname{SO}(V,q)}_{\operatorname{hor}}$, the map
$\mathcal{A}_{Q^{B}}\to\Omega^{2}(Q^{B},V)^{\operatorname{SO}(V,q)}_{\operatorname{hor}},\quad\tilde{\varphi}\mapsto\mathrm{d}^{\tilde{\varphi}}\theta_{\mathcal{B}}$ (48)
@@
Note that the space of spinorial connections $\mathcal{A}^{s}_{Q}$ contains those connections that mimic the Levi-Civita connection along the bosonic distribution. A precise relation will be presented in section 5.
@@
$\mathcal{A}^{s}_{Q}\cong\mathcal{A}_{Q^{\mathbb{F}}}$ (50)